Explain tensile testing for ductile and brittle materials. [14]

2. What is Cast Iron? Explain different types of CI with industrial uses. [7]

3. What is hardness? How do you measure it? [7]

4. Classify engineering materials. [7]

5. Explain following properties [7]


i. Strength

ii. Fatigue

iii. Ductility

iv. Toughness

6. A hollow right circular cylinder is made of certain material and has an outside

diameter of 75mm and an inside diameter of 60mm. The cylinder measures 600

mm in length and is subjected to axial compressive load of 50kN. Determine

i. Stress

ii. Deformation

iii. Material type

Take the modulus of elasticity of the material to be 100 GPa. [8]


1
Expert's answer
2021-01-18T01:12:12-0500

1-

In brittle materials, the UTS will at the end of the linear-elastic portion of the stress-strain curve or close to the elastic limit. In ductile materials, the UTS will be well outside of the elastic portion into the plastic portion of the stress-strain curve

2-It is the combination of high carbon content, and the presence of silicon, that gives cast iron its excellent castability. Various types of cast irons are produced using different heat treatment and processing techniques, including gray iron, white iron, malleable iron, ductile iron, and compacted graphite iron

3-hardness test is typically performed by pressing a specifically dimensioned and loaded object (indenter) into the surface of the material you are testing. The hardness is determined by measuring the depth of indenter penetration or by measuring the size of the impression left by an indenter

4-Basic Classification of Engineering Materials

  • Metals. Metals are polycrystalline bodies which are having the number of differentially oriented fine crystals. ...
  • Non-Metals. ...
  • Difference between Metals and Non-Metals.
  • Other classification of engineering materials:
  • Metals and Alloys.
  • Ceramic Materials.
  • Organic Materials


5- i- Definition. In the mechanics of materials, the strength of a material is its ability to withstand an applied load without failure or plastic deformation. ... The applied loads may be axial (tensile or compressive), or rotational (strength shear).

ii - Material fatigue is a phenomenon where structures fail when subjected to a cyclic load. This type of structural damage occurs even when the experienced stress range is far below the static material strengthFatigue is the most common source behind failures of mechanical structures.

iii- Ductility is the ability of a material to be drawn or plastically deformed without fracture. It is therefore an indication of how 'soft' or malleable the material is.

iv- One definition of material toughness is the amount of energy per unit volume that a material can absorb before rupturing. ... It is also defined as a material's resistance to fracture when stressed. Toughness requires a balance of strength and ductility.

6- "\sigma = \frac{F}{A} = \frac{4F}{\pi(0.075\times0.075-0.06\times0.06)}"


= 31438 kpa

= 31.438 mpa

"\Delta L = \frac{\Delta L}{E}"

= 0.188 mm